# ZeroScene: A Zero-Shot Framework for 3D Scene Generation from a Single Image and Controllable Texture Editing

### v1

@@ 3.3 Texture Editing
To guarantee precise alignment between the generated textures and the underlying geometry, we introduce a geometry-aware module that guides the generation process of the diffusion model. This module enforces texture-to-surface conformity by incorporating spatial constraints and geometric priors. Specifically, we render normal maps, position maps, and depth-based edge detection maps of the input mesh from $N$ predefined camera viewpoints. These 2D information are concatenated into a multi-channel tensor $\mathcal{G}\in\mathbb{R}^{H\times W\times 7}$, which serves as the conditioning input to a ControlNet [54]. Here, the normal map characterizes the surface orientation at each point, ensuring that texture details correspond to the bump variations of the 3D surface. The position map stores the $XYZ$ coordinates of 3D points, acting as spatial anchors to facilitate logical texture distribution across structural partitions. The edge map delineates object boundaries, constrains the valid regions for texture synthesis, and enhances structural articulation along edges. This geometric condition $\mathcal{G}$ is encoded via a branch parallel to the main text-to-image diffusion model backbone [40], and is injected into the denoising process through cross-attention mechanisms to integrate both geometric and textual features, thereby achieving ”texture follows geometry” within a single view. The learning objective is defined as:
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{\mathbf{z},\mathbf{c},y,\bm{\epsilon},t}\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{t};\mathbf{c},y,t)\right\|_{2}^{2}$ (2)
@@
We observe that when generating only based on the conditional image corresponding to the target new view, it often fails to ensure texture consistency of the overlapping regions between the new image and the existing views, which adversely affects the quality of subsequent texture synthesis. Inspired by relative work [36], we propose a mask-guided progressive image generation strategy to enhance global consistency across the multi-view image set, as shown in Fig. 5. Concretely, starting from an initial reference view $v_{1}$, we generate the first complete RGB image $\mathcal{I}_{1}$ using a geometry-aware diffusion model. For each subsequent target view $v_{i}$, we first project all previously generated views $\{\mathcal{I}_{j}\}_{j=1}^{i-1}$ onto the 2D plane of $v_{i}$ according to their camera parameters, resulting in an incomplete view containing known texture information. We then generate a binary mask $\hat{m}$ by identifying regions that are visible in the current view $v_{i}$ but were occluded in all previously generated views. This mask indicates the unknown regions where new texture needs to be synthesized, assigning 0 to these regions and 1 to known areas. Based on this mask, we initialize the latent noise as follows:
@@
where $z_{t}^{\text{proj}}$ denotes the latent representation projected from known regions, and $z_{t}^{\text{rand}}$ is randomly initialized noise. The noise map along with the geometric conditioning image of the current view $v_{i}$, is fed into the diffusion model to perform an inpainting process. This strategy allows the model to generate semantically plausible and visually coherent textures in the masked unknown regions, while maintaining seamless continuity with known areas and strictly adhering to the geometric constraints of the current view. By iterating this ”projection-mask construction-inpainting” pipeline view-by-view, our method facilitates the propagation and expansion of texture information across different views, ultimately yielding a highly consistent multi-view image set $\{\mathcal{I}_{i}\}_{i=1}^{N}$.
Figure 5: Mask-guided progressive image generation strategy. ”Project” refers to the process of projecting all currently known regions into the latent space of the next viewpoint. ”DM” denotes a pre-trained text-to-image diffusion model [40] equipped with ControlNet [54].
@@
After acquiring the RGB image set, it must be mapped into UV space for texture synthesis. Prior to this, the raw images undergo two critical preprocessing steps. First, illumination consistency is addressed using a delighting model [41] to eliminate inter-view illumination discrepancies caused by the native generation of diffusion models, ensuring overall color coherence. Then, each image is super-resolved [43] to a resolution of 2K to enhance texture details. As illustrated in Fig. 6, to back project the images into UV space, a per-pixel confidence map $\mathcal{C}_{i}$ is computed for each view $\mathcal{I}_{i}$, quantifying the reliability of that view contribution to the texture at each point on the model surface. Concretely, we first evaluate the angle between the surface normal and the viewing direction based on the normal map. Surfaces facing the camera directly are assigned higher weights, while pixels with an angle exceeding a predefined threshold $\alpha$ have their confidence set to zero to avoid texture distortion and undersampling caused by excessive grazing angles. To further mitigate potential seams along model contours or depth discontinuous regions, detected edge areas are also excluded from the mask and their confidence is likewise set to zero. This geometric confidence is then multiplied by the prior weight of view $w_{i}\in\mathcal{W}$ to produce the final confidence map and determine the valid pixels available for texture back-projection. Finally, we traverse all valid pixels in the current view, combine them with the UV unwrapping coordinates of the 3D model, and perform weighted mapping of color values onto the UV space based on confidence. After processing all $N$ views, the final texture map $\mathcal{T}$ is obtained by normalizing the accumulated color values with the total confidence. To further enhance rendering realism, we build upon the generated albedo map and employ Hunyuan3D-Paint 2.1 [14, 18] to estimate metallic, roughness, and bump maps, enabling physically based rendering.
Figure 7: Qualitative comparisons between ZeroScene and current state-of-the-art single-image scene generation methods. From left to right: input image, Hunyuan 3D 2.5 [27], MIDI [22], SceneGen [31], and ZeroScene. The images in the first column from top to bottom are sourced from: real photographs (rows 1-5), VLM-generated images (rows 6-7), and 3D-FRONT [15] (row 8).
@@ 4.1 Implementation Details
During the layout optimization stage in Sec. 3.1, the loss weights in Eq. 1 are set as $\lambda_{1}=1$ and $\lambda_{2}=5\times 10^{-2}$. Each instance is trained for 20 epochs, with 2,000 iterations per epoch. For the first 1,200 iterations, only the 3D CD loss is optimized; in the subsequent 800 iterations, the 2D projected CD loss is additionally incorporated for joint optimization. The epoch with the lowest loss value throughout the training process is selected as the final optimized result. The Adam optimizer is employed, with the learning rate for translation, rotation, and scaling parameters set to 0.01. In Sec. 3.3, we predefine $N=10$ camera views, including six principal views (weight $\mathcal{W}=1$) and four oblique views (elevations $\in\{-20^{\circ},20^{\circ}\}$, azimuths $\in\{45^{\circ},135^{\circ},225^{\circ},325^{\circ}\}$, weight $\mathcal{W}=0.1$). The angular threshold is set to $\alpha=60^{\circ}$. All experiments are conducted on a single NVIDIA A100 GPU with 40GB of memory.
[…]
Hunyuan3D 2.5 [27] - 0.0226 - 72.43 - 0.876 - 0.837
MIDI [22] 0.0409 0.0384 42.76 65.58 0.814 0.841 0.782 0.819
SceneGen [31] 0.0223 0.0416 63.63 67.17 0.835 0.827 0.846 0.812
Ours 0.0163 0.0137 79.35 83.21 0.908 0.913 0.886 0.893
@@
Table 2: Quantitative analysis of ablation study. ”Inp.” indicates image inpainting.
@@
4.2 3D Scene Generation
Baseline & Benchmark. We compare ZeroScene with several representative single image 3D scene generation methods, including Hunyuan3D 2.5 [27], MIDI [22], and SceneGen [31]. All comparative methods were implemented using their publicly available source code and pre-trained models. The evaluation was conducted on an image set containing multiple foreground objects with significant occlusion relationships, which comprises real photographs, content generated via VLM [25], and the publicly available indoor synthetic scene benchmark 3D-FRONT [15].
@@
Qualitative comparison. Since the selected baseline methods lack background generation capabilities, for a fair comparison, Fig. 7 presents exclusively on the foreground objects generation results of each approach under the same scene input. In terms of geometry and texture quality, Hunyuan 3D 2.5 may exhibit partial geometry asset loss and texture estimation bias in certain scenarios. When the input image contains objects in close proximity or with inclusion relationships, MIDI tends to generate repetitive structures. SceneGen, which employs Trellis [47] as its 3D backbone, suffers from illumination estimation errors in unseen viewpoints during the generation process, leading to localized texture overexposure or underexposure. Additionally, due to the lack of completion processing for segmented instance images, this method may produce erroneous redundant structures when handling occluded objects. In contrast, our method excels in preserving structural integrity of individual objects and high-frequency texture details, while also enhancing rendering realism through PBR materials. Regarding layout construction, Hunyuan 3D 2.5 natively captures image features and produces spatially reasonable asset distributions. However, the multi-instance attention mechanism adopted by MIDI struggles to accurately model spatial relationships between objects in complex scenes. SceneGen, on the other hand, exhibits noticeable object misalignment. Our method achieves accurate reasoning of spatial relationships between objects and ensures multi-view consistency across the entire scene. Additional qualitative results (including background) generated by ZeroScene can be found in Fig. 12 and 13.
Quantitative comparison. We evaluate the generated 3D scenes from both geometric structure and visual quality perspectives. In terms of geometric structure, we sample discrete points from the surfaces of both the generated assets and the ground truth models (i.e., 3D-FRONT [15]), and align the reconstructed point cloud with the ground truth point cloud. Subsequently, common geometry evaluation metrics, including Chamfer Distance and F-Score, are computed at both the object level and the scene level. For visual quality, we render the full generated scene as well as individual instances based on the camera parameters of the input images, using the input scene images and instance segmentation images as ground truth. CLIP [34] similarity and DINOv2 [32] similarity are adopted as metrics to quantify the texture quality of the generated assets.
As shown in Tab. 1, ZeroScene consistently outperforms existing methods across both geometric and visual dimensions, at both the object and scene levels. This advantage stems from its ability to effectively capture local asset details and accurately recover scene layouts in a joint 3D and 2D space. Moreover, among all compared methods, ZeroScene is the only one that supports separating objects from the scene and specifically processing the background, further enhancing its overall performance and usability.
@@ 4.3 Texture Editing
Baseline & Benchmark. We compare our method against state-of-the-art texture generation approaches, including MVPaint [10], MV-Adapter [23], and EmbodiedGen [44]. To comprehensively evaluate the generalization capability of text-to-texture generation methods across various mesh structures, we construct a diverse and compact benchmark. The models in this benchmark are randomly sampled from 3D datasets (i.e., Objaverse-XL [11], GSO [13]), with additional scene assets generated by ZeroScene.
Qualitative comparison. Fig. 9 presents a comparative visualization of texture synthesis results on the same mesh model under identical textual prompts, generated by various methods. As demonstrated in the results, MVPaint struggles to adequately capture key semantic information from the text and exhibits noticeable inconsistencies across different viewpoints. MV-Adapter employs a multi-view diffusion model to generate textures for textureless meshes; however, its output images often retain illumination interference, leading to visible artifacts after texture mapping. The underlying image generation model used in EmbodiedGen occasionally fails to produce multi-view content that aligns closely with the given text and is prone to distortion during back projection onto the mesh surface. In contrast, our method generates textures with high clarity and rich details that accurately reflect the visual descriptions in the prompt, while also demonstrating excellent multi-view consistency. Furthermore, by incorporating PBR materials, the textured mesh exhibits enhanced three dimensional depth and photorealism. Additional comparative results can be found in Fig. 14 and Fig. 15.
@@
MVPaint [10] 54.77 10.15 0.805
MV-Adpater [23] 51.48 8.34 0.817
EmbodiedGen [44] 63.62 11.97 0.733
@@
Table 4: Quantitative analysis of ablation study. ”Mask-guided.” denotes the mask-guided image generation strategy, and ”PBR Mat.” refers to the estimation of PBR materials.
@@
Quantitative comparison. Upon completing the texture generation for the 3D asset, we uniformly sample 36 viewpoints around the mesh model at a $10^{\circ}$ elevation angle, rendering images at a resolution of $512\times 512$. For models with existing textures from the Objaverse-XL [11] and GSO [13] datasets, we follow the same sampling procedure and use their renderings as ground truth. Subsequently, we compare the distribution of the rendered image set obtained from the generated textures against that of the image set sampled from the real textures. To quantify the distribution discrepancy and visual quality between the two image sets, we adopt widely-used evaluation metrics in text-to-texture generation tasks: Fréchet Inception Distance (FID) [19] and Kernel Inception Distance (KID) [4]. Additionally, we compute the CLIP similarity to further assess the semantic consistency between the generated textures and the input text prompts. As shown in Tab. 3, our method achieves lower FID and KID scores, indicating that the visual distribution of the generated textures is closer to that of real object textures. Meanwhile, the higher CLIP score demonstrates that our approach can produce textures that more accurately align with the user’s input intent.
Figure 10: Ablation study. The prompt is ”a plastic zebra toy”. The complete design enables the generated textures to have superior consistency and rendering quality.

### v2

@@ 3.3 Texture Editing
To guarantee precise alignment between the generated textures and the underlying geometry, we introduce a geometry-aware module that guides the generation process of the diffusion model. This module enforces texture-to-surface conformity by incorporating spatial constraints and geometric priors. Specifically, we render normal maps, position maps, and depth-based edge detection maps of the input mesh from $N$ predefined camera viewpoints. These 2D information are concatenated into a multi-channel tensor $\mathcal{G}\in\mathbb{R}^{H\times W\times 7}$, which serves as the conditioning input to a ControlNet [ZRA23]. Here, the normal map characterizes the surface orientation at each point, ensuring that texture details correspond to the bump variations of the 3D surface. The position map stores the $XYZ$ coordinates of 3D points, acting as spatial anchors to facilitate logical texture distribution across structural partitions. The edge map delineates object boundaries, constrains the valid regions for texture synthesis, and enhances structural articulation along edges. This geometric condition $\mathcal{G}$ is encoded via a branch parallel to the main text-to-image diffusion model backbone [Tea24], and is injected into the denoising process through cross-attention mechanisms to integrate both geometric and textual features, thereby achieving "texture follows geometry" within a single view. The learning objective is defined as:
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{\mathbf{z},\mathbf{c},y,\boldsymbol{\epsilon},t}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t};\mathbf{c},y,t)\right\|_{2}^{2}$ (3)
@@
Figure 6: Qualitative comparisons between ZeroScene and current state-of-the-art single-image scene generation methods. From left to right: input image, Hunyuan 3D 2.5 [LZL25a], MIDI [HGA25], SceneGen [MWZX26], and ZeroScene. The images in the first column from top to bottom are sourced from: real photographs (rows 1-5) and 3D-FRONT [FCG21] (rows 6-8).
We observe that when generating only based on the conditional image corresponding to the target new view, it often fails to ensure texture consistency of the overlapping regions between the new image and the existing views, which adversely affects the quality of subsequent texture synthesis. Inspired by relative work [RMA23], we propose a mask-guided progressive image generation strategy to enhance global consistency across the multi-view image set, as shown in Fig. 4. Concretely, starting from an initial reference view $v_{1}$, we generate the first complete RGB image $\mathcal{I}_{1}$ using a geometry-aware diffusion model. For each subsequent target view $v_{i}$, we first project all previously generated views $\{\mathcal{I}_{j}\}_{j=1}^{i-1}$ onto the 2D plane of $v_{i}$ according to their camera parameters, resulting in an incomplete view containing known texture information. We then generate a binary mask $\hat{m}$ by identifying regions that are visible in the current view $v_{i}$ but were occluded in all previously generated views. This mask indicates the unknown regions where new texture needs to be synthesized, assigning 0 to these regions and 1 to known areas. Based on this mask, we initialize the latent noise as follows:
@@
where $z_{t}^{\text{proj}}$ denotes the latent representation obtained by encoding the projection of known regions via a Variational Autoencoder, and $z_{t}^{\text{rand}}$ is randomly initialized noise. The noise map along with the geometric conditioning image of the current view $v_{i}$, is fed into the diffusion model to perform an inpainting process. This strategy allows the model to generate semantically plausible and visually coherent textures in the masked unknown regions, while maintaining seamless continuity with known areas and strictly adhering to the geometric constraints of the current view. By iterating this "projection-mask construction-inpainting" pipeline view-by-view, our method facilitates the propagation and expansion of texture information across different views, ultimately yielding a highly consistent multi-view image set $\{\mathcal{I}_{i}\}_{i=1}^{N}$.
[…]
Hunyuan3D 2.5 [LZL25a] - 0.0226 - 72.43 - 0.876 - 0.837
MIDI [HGA25] 0.0409 0.0384 42.76 65.58 0.814 0.841 0.782 0.819
SceneGen [MWZX26] 0.0223 0.0416 63.63 67.17 0.835 0.827 0.846 0.812
Ours 0.0163 0.0137 79.35 83.21 0.908 0.913 0.886 0.893
Table 2: Runtime analysis.
Stage
Instance Segmentation
and Generation
Point Cloud
Extraction
Layout
Optimization
Background
Handling
Total
Time 149.32s 21.33s 69.93s 63.17s $\sim$300s
@@
After acquiring the RGB image set, it must be mapped into UV space for texture synthesis. Prior to this, the raw images undergo two critical preprocessing steps. First, illumination consistency is addressed using a delighting model [Tea25] to eliminate inter-view illumination discrepancies caused by the native generation of diffusion models, ensuring overall color coherence. Then, each image is super-resolved [WXDS21] to a resolution of 2K to enhance texture details. As illustrated in Fig. 5, to back project the images into UV space, a per-pixel confidence map $\mathcal{C}_{i}$ is computed for each view $\mathcal{I}_{i}$, quantifying the reliability of that view contribution to the texture at each point on the model surface. Concretely, we first evaluate the angle between the surface normal and the viewing direction based on the normal map. Surfaces facing the camera directly are assigned higher weights, while pixels with an angle exceeding a predefined threshold $\alpha$ have their confidence set to zero to avoid texture distortion and undersampling caused by excessive grazing angles. To further mitigate potential seams along model contours or depth discontinuous regions, detected edge areas are also excluded from the mask and their confidence is likewise set to zero. This geometric confidence is then multiplied by the prior weight of view $w_{i}\in\mathcal{W}$ to produce the final confidence map and determine the valid pixels available for texture back-projection. Finally, we traverse all valid pixels in the current view, combine them with the UV unwrapping coordinates of the 3D model, and perform weighted mapping of color values onto the UV space based on confidence. After processing all $N$ views, the final texture map $\mathcal{T}$ is obtained by normalizing the accumulated color values with the total confidence. To further enhance rendering realism, we build upon the generated albedo map and employ Hunyuan3D-Paint 2.1 [FYY25, HYY25] to estimate metallic, roughness, and bump maps, enabling physically based rendering.
@@ 4.1 Implementation Details
During the layout optimization stage in Sec. 3.1, the loss weights in Eq. 1 are set as $\lambda_{1}=1$ and $\lambda_{2}=5\times 10^{-2}$. Each instance is trained for 20 epochs, with 2,000 iterations per epoch. For the first 1,200 iterations, only the 3D CD loss is optimized; in the subsequent 800 iterations, the 2D projected CD loss is additionally incorporated for joint optimization. The epoch with the lowest loss value throughout the training process is selected as the final optimized result. The Adam optimizer is employed, with the learning rate for translation, rotation, and scaling parameters set to 0.01. The aforementioned parameter configuration exhibits excellent generalization ability when processing images from diverse sources, enabling a stable optimization process and accurate estimation of layout parameters. In Sec. 3.3, we predefine $N=10$ camera views, including six principal views (weight $\mathcal{W}=1$) and four oblique views (elevations $\in\{-20^{\circ},20^{\circ}\}$, azimuths $\in\{45^{\circ},135^{\circ},225^{\circ},325^{\circ}\}$, weight $\mathcal{W}=0.1$). The angular threshold is set to $\alpha=60^{\circ}$. All experiments are conducted on a single NVIDIA A100 GPU with 40GB of memory.
4.2 3D Scene Generation
Baseline & Benchmark. We compare ZeroScene with several representative single image 3D scene generation methods, including Hunyuan3D 2.5 [LZL25a], MIDI [HGA25], and SceneGen [MWZX26]. All comparative methods were implemented using their publicly available source code and pre-trained models. The evaluation was conducted on an image set containing multiple foreground objects with significant occlusion relationships, which comprises real photographs, content generated via VLM [HLG24], and the publicly available indoor synthetic scene benchmark 3D-FRONT [FCG21].
@@
Qualitative comparison. Since the selected baseline methods lack background generation capabilities, for a fair comparison, Fig. 6 presents exclusively on the foreground objects generation results of each approach under the same scene input. In terms of geometry and texture quality, Hunyuan 3D 2.5 may exhibit partial geometry asset loss and texture estimation bias in certain scenarios. When the input image contains objects in close proximity or with inclusion relationships, MIDI tends to generate repetitive structures. SceneGen, which employs Trellis [XLX25] as its 3D backbone, suffers from illumination estimation errors in unseen viewpoints during the generation process, leading to localized texture overexposure or underexposure. Additionally, due to the lack of completion processing for segmented instance images, this method may produce erroneous redundant structures when handling occluded objects. In contrast, our method excels in preserving structural integrity of individual objects and high-frequency texture details, while also enhancing rendering realism through PBR materials. Regarding layout construction, Hunyuan 3D 2.5 natively captures image features and produces spatially reasonable asset distributions. However, the multi-instance attention mechanism adopted by MIDI struggles to accurately model spatial relationships between objects in complex scenes. SceneGen, on the other hand, exhibits noticeable object misalignment. Our method achieves accurate reasoning of spatial relationships between objects and ensures multi-view consistency across the entire scene. Additional qualitative results (including background) generated by ZeroScene can be found in Fig. 11 and 12.
Table 3: Quantitative analysis of ablation study. "Inp." indicates image inpainting.
@@
Quantitative comparison. We evaluate the generated 3D scenes from both geometric structure and visual quality perspectives. In terms of geometric structure, we sample discrete points from the surfaces of both the generated assets and the ground truth models (i.e., 3D-FRONT [FCG21]), and align the reconstructed point cloud with the ground truth point cloud. Subsequently, common geometry evaluation metrics, including Chamfer Distance and F-Score, are computed at both the object level and the scene level. For visual quality, we render the full generated scene as well as individual instances based on the camera parameters of the input images, using the input scene images and instance segmentation images as ground truth. CLIP [RKH21] similarity and DINOv2 [ODM24] similarity are adopted as metrics to quantify the texture quality of the generated assets.
As shown in Tab. 1, ZeroScene consistently outperforms existing methods across both geometric and visual dimensions, at both the object and scene levels. This advantage stems from its ability to effectively capture local asset details and accurately recover scene layouts in a joint 3D and 2D space. Moreover, among all compared methods, ZeroScene is the only one that supports separating objects from the scene and specifically processing the background, further enhancing its overall performance and usability. Table 2 provides an analysis of the average runtime of ZeroScene for 3D scene generation from a single image.
@@ 4.3 Texture Editing
Baseline & Benchmark. We compare our method against state-of-the-art texture generation approaches, including TEXTure [RMA23], MVPaint [CMZ25], MV-Adapter [HGW25], and EmbodiedGen [WLC25]. To comprehensively evaluate the generalization capability of text-to-texture generation methods across various mesh structures, we construct a diverse and compact benchmark. The models in this benchmark are randomly sampled from 3D datasets (i.e., Objaverse-XL [DLW23], GSO [DFK22]), with additional scene assets generated by ZeroScene.
Qualitative comparison. Fig. 8 presents a comparative visualization of texture synthesis results on the same mesh model under identical textual prompts, generated by various methods. As demonstrated in the results, TEXTure tends to generate noticeable seam artifacts and texture stretching at view transition regions when handling complex geometric topologies. MVPaint struggles to adequately capture key semantic information from the text and exhibits noticeable inconsistencies across different viewpoints. MV-Adapter employs a multi-view diffusion model to generate textures for textureless meshes; however, its output images often retain illumination interference, leading to visible artifacts after texture mapping. The underlying image generation model used in EmbodiedGen occasionally fails to produce multi-view content that aligns closely with the given text and is prone to distortion during back projection onto the mesh surface. In contrast, our method generates textures with high clarity and rich details that accurately reflect the visual descriptions in the prompt, while also demonstrating excellent multi-view consistency. Furthermore, by incorporating PBR materials, the textured mesh exhibits enhanced three dimensional depth and photorealism. Additional comparative results can be found in Fig. 13 and Fig. 14.
@@
TEXTure [RMA23] 57.31 9.76 0.811
MVPaint [CMZ25] 54.77 10.15 0.805
MV-Adpater [HGW25] 51.48 8.34 0.817
EmbodiedGen [WLC25] 63.62 11.97 0.733
@@
Quantitative comparison. Upon completing the texture generation for the 3D asset, we uniformly sample 36 viewpoints around the mesh model at a $10^{\circ}$ elevation angle, rendering images at a resolution of $512\times 512$. For models with existing textures from the Objaverse-XL [DLW23] and GSO [DFK22] datasets, we follow the same sampling procedure and use their renderings as ground truth. Subsequently, we compare the distribution of the rendered image set obtained from the generated textures against that of the image set sampled from the real textures. To quantify the distribution discrepancy and visual quality between the two image sets, we adopt widely-used evaluation metrics in text-to-texture generation tasks: Fréchet Inception Distance (FID) [HRU17] and Kernel Inception Distance (KID) [BSAG18]. Additionally, we compute the CLIP similarity to further assess the semantic consistency between the generated textures and the input text prompts. As shown in Tab. 4, our method achieves lower FID and KID scores, indicating that the visual distribution of the generated textures is closer to that of real object textures. Meanwhile, the higher CLIP score demonstrates that our approach can produce textures that more accurately align with the user’s input intent.
Table 5: Quantitative analysis of ablation study. "Mask-guided." denotes the mask-guided image generation strategy, and "PBR Mat." refers to the estimation of PBR materials.
@@
Figure 9: Ablation study. The prompt is "a plastic zebra toy". The complete design enables the generated textures to have superior consistency and rendering quality.